- Press Release
- Oct 2, 2022
NOAA GOES SATOPS Morning Report 7 Feb 2001
SATOPS MORNING REPORT February 7, 2001
Senior Duty Meteorologist advised at 2130z on February 6 that a Critical
Weather Days have been declared for 1200z February 7 until 1200z February
No change in the status of the POES spacecraft.
* Frame Synchronizer testing continues.
NOAA-14 rev 31470 / F at 1556z on February 6: L5B was noisy from spacecraft
with 97% of the expected data received. All L5B recovered cleanly
on rev 31472.
NOAA-14 rev 31480 / W at 0907z on February 7: L5B was noisy from spacecraft
with 81.4% of the expected data received. L5B not recoverable due
to spacecraft recorder sequence constraints.
No change in the status of the DMSP spacecraft.
OPS47 rev 30315 / POGO at 0639z on February 7: All data recovered
clean at SOCC, but Air Force Weather Agency had ground system ingest problems
with the data.
No change in the status of the GOES spacecraft.
GOES operations were nominal over the past 24 hours.
Questions or comments to firstname.lastname@example.org